OpenStudy (anonymous):

Help: If a circle has a circumference of C, what is the area of the circle in terms of C? A=πC A=C2/4π A=2C/π A=2C2

OpenStudy (anonymous):

i dnt know i have a feeling its C but im not quite shure @Rushwr

OpenStudy (jhannybean):

\[\sf C=2\pi r\]\[\sf A=\pi r^2\]If we use our equation of circumference and solve for r, we wcan input that into our equation for Area to set it in terms of C \[\sf C = 2\pi r \implies r = \frac{C}{2\pi}\]\[\sf A= \pi\left(\frac{C}{2\pi}\right)^2\]

OpenStudy (jhannybean):

Now let's expand our function of Area. \[\sf A =\cancel{ \pi} \left(\frac{C^2}{4 \cancel{\pi^2}\pi}\right)\]\[\sf \boxed{A = \frac{C^2}{4\pi}}\]

OpenStudy (jhannybean):

Do you follow, @bruno101 ?

OpenStudy (anonymous):

yes i see it seems a little complicated but once u follow its easier @Jhannybean

OpenStudy (jhannybean):

When they say `in terms of C` they want the function (Area) in terms of Circumference. So basically... \(A(\text{circumference})\)

OpenStudy (anonymous):

OK I SEE THANKS!! @Jhannybean

OpenStudy (jhannybean):

No problem :)